OpenStudy (anonymous):

10.)Choose the missing step in the given solution to the inequality -x - 10 > 14 + 2x. ------------------------------------ [step one: -x - 10 > 14 + 2x [step two:-3x -10 > 14 [step three: ?? [step four: x < -8 ------------------------------------- a. -3x > 24 b. -3x < 4 c. -3x > -4 d. 8 < x 11.) Choose the correct description of the graph of the inequality x - 12 <= 5x. a. Open circle on -3, shading to the left. b. Closed circle on -3, shading to the left. c. Open circle on -3, shading to the right. d. Closed circle on -3, shading to the right.

OpenStudy (anonymous):

10) would be a) as you take add 10 on each side of the inequality.

OpenStudy (anonymous):

13.) dentify the slope and y-intercept of the inequality 8x + 2y > 12. a. slope = 4, y-intercept = 6 b. slope = -4, y-intercept = 12 c. slope = -4, y-intercept = 6 d. slope = -1/4, y-intercept = 12 14.) Solve the inequality 6x - 4 > 2x + 12 a. x > 4 b. x > 1 c. x < 1 d. x < 4

OpenStudy (anonymous):

15.) Choose the correct description of the graph of the compound inequality x - 2 > -4 and 3x <= 15 a. A number line with an open circle on -2, shading to the left, and a closed circle on 5, shading to the right. b. A number line with a closed circle on -2, shading to the left, and an open circle on 5, shading to the right. c. A number line with an open circle on -2, a closed circle on 5, and shading in between. d. A number line with a closed circle on -2, an open circle on 5, and shading in between.

OpenStudy (anonymous):

I need help with those as well D;

OpenStudy (anonymous):

please help? D: